OpenStudy (anonymous):

The total cost for 8 bracelets, including shipping was $54. The shipping charge was $6. Define your variable and write an equation that models the cost of each bracelet. Is this wrong or right? x8 y54 -$6 -8 -8 54 -14 = 40 40/8 = 5

OpenStudy (anonymous):

i think the answer is $6 for each bracelet

OpenStudy (taramgrant0543664):

I'm unsure of what you are doing but I'll show you how I would do it. Since there is only one shipping charge we don't need to give it a variable. Together the shipping ant the eight bracelets equal 56. I will let x represent the price of one bracelet and since there is 8 I will multiply x by 8. So my equation will be: 8x+6=56 8x=50 x=6.25

OpenStudy (anonymous):

54-6=48 48/8=6

OpenStudy (taramgrant0543664):

oppps just realized I put 56 not 54v sorry!! So 8x+6=54 8x=48 x=6

OpenStudy (anonymous):

But didn't they say define my variables? Don't i need to make 6 y or something?